Robinson Industries has a defined benefit pension plan that specifies annual retirement benefits equal to:

1.9% x Service years x Final Year’s salary

Patty Mills was hired by Robinson at the beginning of 2002. Mills is expected to retire at the end of 2041 after 40 years of service. His retirement is expected to span 20 years. At the end of 2016, 15 years after being hired, his salary is $65,500. The company's actuary projects Mills' salary to be $86,700 at retirement. The actuary's discount rate is 8%.

PVA Factors

PVA, n=15, i=8%      8.55948

PVA, n=20, i=8%      9.81815

PVA, n=25, I -8%      10.67478

PV Factors

PV, n=15, i=8%        .31524

PV, n=20, i=8%        .21455

PV, n=25. I =8%       .14602

What is the company's projected benefit obligation at the end of 2016 with respect to Patty Mills? ___________________
Answer #1

TOTAL ANNUAL BENEFITS ON RETIREMENT 1.9%×SERVICE YEARS X FINAL YEAR SALARY

ANNUAL BENEFIT =0.019×40×86700=65892

Now retirement last 20 years

PVA n=20 r=8% is PVA=9.81815

SO TOTAL VALUE OF FUND =65892×9.81815

RETIREMENT FUND=646937.54

After 15 years means 25 before retirement

PV FACTOR FOR 8% AND 25 YEAR

PV=0.14602

VALUE OF BENEFIT =PV X RETIREMENT FUND

VALUE OF BENEFIT =0.14602×646937.54

VALUE OF BENEFIT =94465.82

At end of 2016 company projected BENEFIT obligation with respect to party mills is 94465.82$ which is required answer